# Recent Advances in Discrete Speech Tokens: A Review

### v1

@@ VI-B3 Non-Autoregressive TTS
While autoregressive modeling is the current mainstream of TTS with discrete tokens, non-autoregressive models also exist. A convenient way is to treat the code-vectors of discrete tokens as continuous features and use the well-established regression objective to model these vectors. An example is NaturalSpeech 2 [248] which uses diffusion model to generate the sum of SoundStream code-vectors per frame, with duration-upsampled phone representations as conditions. As tokens are inherently discrete, discrete diffusion models [249] are also applied. UniCATS [119] and NaturalSpeech 3 [27] use discrete diffusion to model semantic and disentangled acoustic tokens respectively. Similarly, the masked prediction of tokens proposed by MaskGIT [51] is also incorporated into TTS [143, 250]. These non-autoregressive methods are naturally more robust than autoregressive methods in inference, and also supports speech editing.
@@ VI-B4 Unified Speech Generation
The language modeling approach of discrete tokens allows a unified generation framework for multiple tasks. It suffices to use a task identifier to condition the unified language model. VALL-E X [251] uses VALL-E to combine cross-lingual TTS and speech-to-speech translation. Make-A-Voice [252] uses a hierarchical Transformer architecture, and supports TTS, VC and SVS and singing VC. SpeechX [253] includes more tasks such as speech editing and noise suppression with an AR+NAR codec language model like VALL-E. UniAudio [242] designs a hierarchical Transformer to form a single-stage autoregressive generation process of acoustic tokens only. UniAudio supports 11 speech and audio generation tasks within a single model. It achieves comparable or even better performance than special-purpose in most tasks, demonstrating the potential of a large-scale foundation model.
@@ VI-B5 Limitations
In addition to discrete tokens, another rapidly emerging framework for speech generation is diffusion or flow matching-based models. Notable examples include VoiceBox [254], HierSpeech++ [255], E1-TTS [256] and F5-TTS [257]. These models typically generate continuous features in a non-autoregressive way, and some even eliminate the need for forced alignments. Owing to the strong capability of diffusion and flow matching algorithms, they also have remarkable generation quality, diversity and controllability. Compared to this paradigm, autoregressive discrete token-based speech generation models usually fall short in terms of generation robustness and, in many cases, efficiency. Additionally, diffusion-based autoregressive speech generation using continuous features is an emerging area of interest [258, 259, 260, 261]. These approaches with continuous features have a higher upper bound for speech quality and intelligibility, as they inherently avoid quantization errors. Therefore, there is an ongoing debate between discrete and continuous representations for speech generation.
@@ VI-C1 Motivation
End-to-end speech generation is one of the most critical goals in the field of speech. Discrete tokens are a core component of existing spoken language models, as they enable the language modeling technique to be applied directly on speech. The models discussed in this subsection are text-free spoken language models (TF-SLMs). We anticipate that a well-trained TF-SLM will be capable of autoregressively generating semantically coherent speech without the need for text transcription guidance.
@@ VI-C2 Existing Efforts
Ever since GSLM [4] and AudioLM [5] proposed the vision of TF-SLMs, building such models remains a significant challenge as of today. This difficulty primarily arises from the lower language modeling efficiency of speech token sequences compared to text, due to their lower semantic information density, longer sequence lengths, and the presence of paralinguistic information [262]. Current advancements in TF-SLMs mainly focus on two strategies: (1) reducing token frame rates and (2) aligning speech with text.
The first approach aims to shorten speech sequences and enhance semantic density by lowering frame rates to 25Hz [4, 138], 20Hz [168] or even $\approx$5Hz [44, 177]. While mitigating sequence length issues to different degrees, they still encounter scalability limitations [263] and compromise reconstruction quality. The second strategy involves aligning speech with text through methods like initializing pre-training with text LLMs [138], reinforcement learning using ASR and text LLM feedback [264], text-speech token interleaving [265], adopting novel architectures applied in text language modeling [266], etc [267, 268]. However, despite incorporating text-related supervisory signals or training techniques during training, these models still struggle to generate semantically reliable long speech during inference due to the lack of explicit transcription guidance. Meanwhile, full duplex modeling has been proposed [269] to enable users to interrupt the generation process and start new dialogue turns at will.
@@ VI-C3 Limitations
Although these methods show promise, achieving semantic coherence is still a challenging goal, leaving significant progress to be made toward the goal of truly end-to-end speech generation. Improving the semantic density and expressiveness of discrete speech representations, making it easier to align text and speech during TF-SLM training, is a promising direction for future exploration.
@@ VI-D1 Motivation
Since TF-SLM remains an open problem, the prevalent successful speech dialogue systems settle for an alternative choice that uses text as explicit guidance. Recent researches, especially following work like OpenAI’s GPT-4o [^13] https://openai.com/index/hello-gpt-4o/, have focused on spoken language models that combine three key capabilities: strong understanding of speech semantics, high-quality speech output, and low latency [270, 271, 272, 10, 273, 164, 274, 275, 276, 277, 278, 163, 279, 280, 281, 282]. We refer to them as text-guided spoken language models (TG-SLMs). Unlike TF-SLMs, while TG-SLMs utilize a unified LLM for seamless processing of user’s speech input and system’s speech output, they internally decompose the end-to-end speech dialogue process into two well-established sub-procedures: spoken language understanding powered by LLMs (SLU), and real-time speech generation (TTS). The two sub-procedures are connected via text as an intermediary to stabilize the semantic coherency of the final output. The LLM first generates a textual response to the audio input, then synthesizes the speech token sequence of the generated text in a streaming fashion. In a TG-SLM, the SLU sub-procedure usually uses continuous speech features as input since they preserve more acoustic details that facilitate understanding, while the TTS sub-procedure typically uses discrete speech tokens as output to better fit the nature of LLM autoregressive generation.
@@ VI-D2 Speech Generation in TG-SLMs
To reduce modeling complexity and better align with the autoregressive generation paradigm of LLMs, TG-SLMs favor single-layer discrete speech tokens as direct LLM outputs. SpeechGPT-Gen [270] applies a chain-of-modality approach, which utilizes the tokens of the first RVQ layer of SpeechTokenizer [9] as the output of LLM (i.e., TTS sub-procedure output), and uses a separate flow matching model to infer the rest RVQ layers after finishing generating the first RVQ layer. GLM-4-Voice [163] uses 12.5Hz supervised semantic tokens as both the speech input and speech output of the LLM. To lower latency and preserve the intelligence of the LLM, it employs an interleaving manner to generate text and corresponding speech token sequences in turn. VITA-1.5 [280] leverages TiCodec [21] as the output tokens and employs a two-stage decoding process: a non-autoregressive decoder for initial semantic processing, followed by an autoregressive decoder for refined speech token generation. OpenOmni [281] deploys the $\mathcal{S}^{3}$ Tokenizer [40] as output and adds a streaming speech decoder to simultaneously generate text and speech responses.
To better rebuild the speech information with the help of pretrained LLMs, several TG-SLM works use multi-layer speech tokens as LLM output. Moshi [10] uses its own Mimi codec with 8 RVQ layers. The text response and Mimi’s first RVQ layer (semantic layer) are generated in parallel through separate output heads. After a brief delay, the remaining layers (acoustic layers) begin parallel generation. Mini-Onmi [272] and Mini-Omni2 [273] utilize multi-layer multi-resolution codec SNAC [22] as the speech output. Different from Moshi, they decode text tokens and each layer of speech tokens in different output heads in parallel to reduce latency. They further employ a parallel decoding strategy similar to MusicGen [243] for fast and high-quality multi-layer speech token generation.
Mainstream TG-SLMs which use discrete tokens as LLM outputs need an additional decoder (whether the decoder of the employed codec [270, 272, 10, 273, 280] or a separately trained one [164, 277, 163, 281, 282]) to synthesize continuous speech signals. Recently, SALMONN-Omni [274] and MinMo [282] propose to streamingly synthesize speech signals directly based on the LLM hidden embedding, eliminating the need for discrete tokens, additional decoders, or even explicit text tokens, hence further improving the real-time ability.
@@ VII Challenges and Future Directions
As highlighted in the preceding sections of this review, discrete speech tokens have emerged as a thriving and central research direction in the era of speech LLMs. However, the current types of discrete speech tokens still exhibit certain limitations and challenges that need to be addressed. In this section, we summarize the existing challenges in this field and outline the corresponding future directions.
@@ VII-1 Low-Bitrate Tokens
Not only is reducing the bitrate desirable for compression and transmission, but it also eases the downstream modeling burden. For bitrates, factors $Q$ (number of quantizers) and $F$ (frame rate) play a more important role than $V$ (vocabulary size). For $Q$, there have been efforts to construct single-codebook acoustic tokens [71, 76, 77, 19] and lower the bitrate to $\leq 0.3$kbps [71, 17, 28]. Using only a single codebook is very beneficial for downstream models since speech can be truly regarded as another “natural language” then. However, there is usually a noticeable degradation in reconstruction performance in these tokens. A critical problem lies in how to better utilize the highly-compact discrete VQ representation space. For $F$, the frame rates of most tokens are still much greater than text sequences, even if some attempt to lower the frame rate down to $\leq 12.5$Hz [10, 161]. As pointed out in [262], the difference in sequence lengths between phoneme and speech tokens significantly influences the syntactic and semantic modeling capabilities of language models trained on these two modalities. It is thus important to further shorten the sequence length of speech tokens. However, reducing the frame rate of tokens may also lead to decreased intelligibility in reconstructed speech, a challenge that needs to be better addressed. As for the vocabulary size $V$, it will not significantly impact the bitrate or downstream modeling generally, provided that it is not excessively large.
For acoustic tokens, it remains an open problem what the lower bound of bitrate and the frame rate $F$ are, and how to reach them. More powerful network architectures or advanced VQ schemes should be helpful, and reducing the temporal redundancy by disentangling global information is also a possible solution.
@@ VII-2 Streaming Ability and Efficiency
Real-time applications require tokens to be streamable. Streaming ability means encoding or decoding tokens in real time, or with minimal latency, once sufficient data is received. This requirement means the token encoder and decoder must generate outputs using only prior inputs or within a fixed chunk size. For most acoustic tokens using CNN blocks, achieving this is easy due to their fixed receptive field. For acoustic tokens with Transformer blocks, an attention mask is essential. However, most SSL models employ a non-causal Transformer architecture, which renders most semantic tokens derived from these models unsuitable for real-time tokenization. Given that these tokens are frequently designed as targets for language models, the token vocoder for them should also support streaming [162]. It remains unclear how much performance degradation would result from transitioning to non-causal architectures in both SSL models and token vocoders.
@@ VII-3 Disentanglement in Acoustic Tokens
Whether disentanglement should be incorporated into acoustic tokens depends on the specific application. If reconstruction is the major objective, disentangling speaker or prosody information from acoustic tokens may not be necessary. However, disentanglement can help reduce the bitrate in time-varying tokens and ensure anonymity during transmission, which is particularly important when privacy is a concern. For downstream modeling tasks, factorizing the information encoded into tokens is desirable for reducing modeling complexity and achieving independent control of different voice properties in generative tasks. There are currently only limited efforts on decoupled acoustic tokens, and the decoupling effect is still suboptimal, e.g. VC performance is lower than state-of-the-art VC models, or a negative impact on reconstruction quality is witnessed [28]. Future work should consider more advanced techniques for information decoupling.
@@ VII-4 Variable Frame Rate Tokens
Current speech tokens are usually designed at a fixed temporal rate, while the underlying linguistic units in speech are highly variable-rate. This can be an important insight, as discrete tokens will have a better alignment and relation with texts then. This offers the potential of further reducing the bitrate, and more importantly, helping downstream tasks utilize the inductive bias in tokens that more resembles units in a natural language. So far, only semantic tokens have explored varying temporal rates, but with a complicated algorithm or iterative training process [44, 177]. More efforts need to be taken on whether it is possible for acoustic tokens to vary the frame rate, and what benefit in practice will these variable frame rate tokens bring.
@@ VII-5 Combining Acoustic and Semantic Tokens
It is widely acknowledged, and also shown in this review, that acoustic tokens emphasize preserving the acoustic details of speech, whereas semantic tokens often discard them. A natural question arises: Does there exist a representation space that is both acoustic and semantic? In other words, can a representation space contain rich speech understanding capabilities while also reconstructing acoustic details at a decent level? It has been proven that incorporating semantic information from SSL models enhances the reconstruction and downstream modeling performance of acoustic tokens [9, 25, 17]. This approach shows promise and should be further optimized to achieve lower bitrates and even a single codebook.
@@ VII-6 Paralinguistics in Semantic Tokens
Although speaker information is generally considered irrelevant for semantic content, prosody serves as a crucial component of paralinguistic information. Semantic tokens derived through simple clustering methods are likely to discard both speaker information and prosody, which can be detrimental to downstream models’ ability to perceive and generate rich emotions, tones, singing voices, and non-verbal vocalizations that convey semantic meaning. This problem can be partially mitigated by certain quantization approaches that encode more information from SSL features [142, 140, 141]. However, these methods typically require more codebooks and a higher bitrate. Supervised tokenization could also be considered for directly guiding tokens toward paralinguistic information. However, current supervised tokenization methods rely solely on ASR, which is insufficient for preserving prosody.
@@ VII-8 Timbre Control in Token Vocoders
For speaker-decoupled acoustic tokens and semantic tokens, the token vocoder should be responsible for controlling speaker timbre. Currently, both end-to-end GAN-based token-to-wav models [183] and flow matching-based token-to-mel models [40] have demonstrated strong timbre control capabilities. It remains an open question whether the upper bound of the former method can be improved by training on large-scale datasets, as is done with the latter. Also, the timbre controllability of in-the-wild reference prompts with various acoustic conditions should be further investigated.
@@ VII-9 Adaptivity
The application scenarios of discrete speech tokens are diverse, ranging from transmission and compression to various speech processing tasks and speech LLMs. It is challenging to develop a single speech tokenizer that can meet all the distinct requirements of these diverse scenarios. However, adaptivity remains a desirable property, where the tokenizer is expected to adapt to different scenarios given a task prompt or a few examples of data. For example, an adaptive tokenizer is allowed to underperform on speech with unseen accents, languages, or acoustic backgrounds, but it should have the capability to quickly adapt to these new scenarios. Another example is that an acoustic tokenizer can determine whether to preserve or remove speaker timbre, or adjust the temporal rate and granularity, based on an external control signal. If such tokens exist and the aforementioned challenges are well addressed, then we may be close to achieving a universal speech tokenizer.
@@ VIII Conclusion
Recently, discrete speech tokens have emerged as a rapidly evolving field and a core research direction in the speech LLM era. These tokens encode acoustic or semantic information into a compact discrete representation space, catalyzing the fusion of LLMs and speech processing, particularly speech generation and spoken language modeling. Existing discrete speech tokens show rich diversity in model architecture and optimization objectives. In this review, we provide a comprehensive introduction to representative categories of discrete speech tokens, summarizing their motivations and limitations. We conduct a unified analysis of reconstruction and voice conversion across different token types to highlight their unique characteristics. We also review efforts to apply discrete tokens to speech processing tasks, including spoken language understanding, speech generation, and spoken language modeling. Finally, we explore future directions for discrete speech tokenization methods. Despite significant progress, substantial development remains ahead. We hope this review lays a solid foundation for future research in speech technology.

### v2

@@ VI-B3 Non-Autoregressive TTS
While autoregressive modeling is the current mainstream of TTS with discrete tokens, non-autoregressive models also exist. These models either treat the code-vectors as continuous features [238], or directly generate discrete tokens by masked prediction [136] or discrete diffusion models [111, 28]. These non-autoregressive methods are naturally more robust than autoregressive methods in inference, and also supports speech editing.
@@ VI-B4 Unified Speech Generation
The language modeling approach of discrete tokens allows a unified generation framework for multiple tasks. It suffices to use a task identifier to condition the unified language model. For example, [239, 240] extends VALL-E with more tasks like cross-lingual TTS, S2ST, speech editing, etc. UniAudio [231] supports 11 speech and audio generation tasks within a single hierarchical Transformer model. Prompt tuning upon a spoken language model has also been explored in [192] for efficient, transferable and versatile generation. These efforts demonstrate the potential of a large-scale foundation model for generation.
@@ VI-B5 Limitations
In contrast to discrete tokens, another emerging framework for speech generation is diffusion or flow matching-based models, including non-autoregressive models [241, 242, 243, 244] or autoregressive models [245, 246, 247, 248, 249]. They generate continuous features, and some even eliminate the need for forced alignments in non-autoregressive generation. Owing to the strong capability of diffusion and flow matching algorithms, they also have remarkable generation fidelity, diversity and controllability. They can have a higher upper bound for speech quality and intelligibility, as they inherently avoid quantization errors. In comparison, discrete token-based speech generation models usually fall short in generation robustness. Therefore, there is an ongoing debate between discrete and continuous representations for speech generation.
@@ VI-C1 Motivation
End-to-end spoken language and dialogue modeling is one of the most ultimate goals in speech technology. Discrete tokens are a core component of existing spoken language models, as they enable the language modeling technique to be applied directly on speech. The models discussed in this subsection are text-free spoken language models (TF-SLMs). We anticipate that a well-trained TF-SLM will be capable of generating semantically coherent speech without the need for text transcription guidance.
@@ VI-C2 Existing Efforts
Ever since GSLM [4] and AudioLM [5] proposed the vision of TF-SLMs, building such models remains a significant challenge even till today. This difficulty primarily arises from the lower language modeling efficiency of speech token sequences compared to text, due to their lower semantic information density, longer sequence lengths, and the presence of paralinguistic information [250]. Current advancements in TF-SLMs mainly focus on two strategies: (1) reducing token frame rates, and (2) aligning speech with text.
The first approach aims to shorten speech sequences and enhance semantic density by lowering frame rates [4, 131, 159] to even $\approx$5Hz [45, 168]. While mitigating sequence length issues to different degrees, they still encounter scalability limitations [251] and compromise reconstruction quality. The second strategy involves aligning speech with text through methods like initializing pre-training with text LLMs [131], reinforcement learning using ASR and text LLM feedback [252], text-speech token interleaving [253], adopting novel architectures applied in text language modeling [254], etc [255, 256]. Meanwhile, full duplex modeling has been proposed [257] to allow users to interrupt and start new dialogues at will. However, despite many efforts, these models still struggle to generate semantically reliable long speech during inference due to the lack of explicit transcription guidance.
@@ VI-C3 Limitations
Although these methods show promise, achieving semantic coherence is still a challenging goal, leaving significant progress to be made toward the goal of truly end-to-end spoken language modeling. Improving the semantic density and expressiveness of discrete speech representations, making it easier to align text and speech during TF-SLM training, is a promising direction for future exploration.
@@ VI-D1 Motivation
Since TF-SLM remains an open problem, the prevalent successful speech dialogue systems settle for an alternative choice that uses text as explicit guidance. Recent researches, especially following work like OpenAI’s GPT-4o [^10] https://openai.com/index/hello-gpt-4o/, have focused on SLMs that combine three key capabilities: strong understanding of speech semantics, high-quality speech output, and low latency [258, 259, 260, 10, 261, 155, 262, 263, 264, 265, 266, 154, 267, 268, 269, 270]. We refer to them as text-guided spoken language models (TG-SLMs). Unlike TF-SLMs, while TG-SLMs utilize a unified LLM for seamless processing of user’s speech input and system’s speech output, they internally decompose the end-to-end speech dialogue process into two well-established sub-procedures: SLU powered by LLMs, and real-time TTS. The two sub-procedures are connected via text as an intermediary to stabilize the semantic coherency of the final output. The LLM first generates a textual response to the audio input, then synthesizes the speech token sequence in a streaming fashion. In a TG-SLM, the SLU sub-procedure usually uses continuous speech features as input since they preserve more acoustic details for understanding, while the TTS sub-procedure typically uses discrete speech tokens as output to better fit LLM autoregressive generation.
@@ VI-D2 Speech Generation in TG-SLMs
To reduce modeling complexity and better align with the autoregressive generation paradigm of LLMs, TG-SLMs favor single-layer discrete speech tokens as direct LLM outputs. Existing works make use of either the first layer of an RVQ codec [258], single-codebook codec [268], or single-codebook supervised semantic token [154, 269]. Specific designs are introduced corresponding to the tokens, such as chain-of-modality [258], token interleaving to lower latency [154], two-stage decoding process [268], etc. To better rebuild the speech information with the help of pretrained LLMs, several TG-SLMs use multi-layer speech tokens as LLM output, such as [10, 260, 261]. They often employ different techniques to generate the text tokens and multi-layer speech tokens in parallel reduce latency.
Mainstream TG-SLMs with discrete tokens as LLM outputs need an additional decoder to synthesize continuous speech signals, either using the codec decoder or a separately-trained vocoder. There are also efforts to streamingly synthesize speech signals directly based on the LLM hidden embedding [262, 270], eliminating the need for discrete tokens, additional decoders, or even explicit text tokens, hence further improving the real-time ability.
@@ VII Challenges and Future Directions
Current discrete speech tokens still exhibit certain limitations and challenges that need to be addressed. In this section, we summarize the existing challenges in this field and outline the corresponding future directions.
@@ VII-1 Low-Bitrate Tokens
For bitrates of tokens, factors $Q$ (number of quantizers) and $F$ (frame rate) play a more important role than $V$ (vocabulary size). Using only a single codebook is very beneficial for language modeling, since speech can be truly regarded as another “natural language” then. However, there is usually a noticeable degradation in reconstruction performance in these tokens. A critical problem lies in how to better utilize the highly-compact discrete VQ representation space. For $F$, the frame rates of most tokens are still much greater than text sequences, which can significantly influence the syntactic and semantic modeling capability of language models [250]. However, reducing the frame rate of tokens may also lead to decreased intelligibility in reconstructed speech. A lower $V$ is also desirable for language modeling and length reduction by BPE.
For acoustic tokens, it remains an open problem what the lower bound of bitrate and the frame rate $F$ are, and how to reach them. More powerful network architectures or advanced VQ strategies should be helpful, and reducing temporal redundancy by disentangling global information is also a promising solution.
@@ VII-2 Streaming Ability and Efficiency
Real-time applications require tokens to be stream-able both in encoding and decoding. For most CNN-based acoustic tokens, achieving this is easy due to their fixed receptive fields. For acoustic tokens with Transformer blocks, an attention mask is necessary. However, most SSL models employ a non-causal Transformer architecture, which makes semantic tokens derived from these models unsuitable for real-time tokenization. It remains unclear how much performance degradation would result from transitioning to causal architectures in both SSL models and token vocoders.
@@ VII-3 Disentanglement in Acoustic Tokens
Whether disentanglement should be incorporated into acoustic tokens depends on the specific application. If reconstruction is the major objective, disentanglement may not be necessary. However, disentanglement can help reduce the bitrate in time-varying tokens, ensure anonymity during transmission, reduce downstream modeling complexity, and achieve independent control of difference voice properties. There are currently only limited efforts on decoupled acoustic tokens, and the decoupling effect is still suboptimal or causing a negative impact on reconstruction quality. More advanced techniques for information decoupling should be considered in the future.
@@ VII-4 Variable Frame Rate Tokens
Current speech tokens are usually designed at a fixed temporal rate, while the underlying linguistic units in speech are highly variable-rate. This can offer an important insight and potential of further reducing the bitrate, and more importantly, closing the gap between speech tokens and natural language units for downstream tasks. So far, only semantic tokens have explored varying temporal rates, but with a complicated algorithm or iterative training process [45, 168]. More explorations need to be taken on variable frame rate acoustic tokens and the benefit of these variable frame rate tokens in practice.
@@ VII-5 Combining Acoustic and Semantic Tokens
Given the distinct properties of acoustic and semantic tokens, a natural question arises: Can a representation space contain rich speech understanding capabilities while also reconstructing acoustic details at a decent level? It is promising to incorporating semantic information from SSL models to enhance the reconstruction and downstream modeling performance of acoustic tokens [9, 26, 18]. This approach should be further optimized to achieve lower bitrates and even a single codebook.
@@ VII-6 Paralinguistics in Semantic Tokens
While speaker information is generally considered irrelevant for semantic content, prosody serves as a crucial component of paralinguistic information. Semantic tokens derived through simple clustering methods are likely to discard both speaker information and prosody, harming downstream models’ ability to handle rich emotions, tones, singing voices, and non-verbal vocalizations that convey semantic meaning. This problem can be partially mitigated by certain VQ approaches that encode more information from SSL features [135, 133, 134], but at a cost of more codebooks and higher bitrates. Supervised tokenization could also be considered for directly guiding tokens toward paralinguistic information. However, current supervised tokens rely solely on ASR, which is insufficient for preserving prosody.
@@ VII-8 Timbre Control in Token Vocoders
For semantic tokens and speaker-decoupled acoustic tokens, token vocoders should be responsible for controlling speaker timbre. Currently, both GAN-based token-to-wav vocoders [171] and flow matching-based token-to-mel models [41] have demonstrated strong timbre control capabilities. It remains an open question whether the upper bound of the former method can be improved by training on large-scale datasets, as is done with the latter. Also, the timbre controllability of in-the-wild reference prompts with various acoustic conditions should be further investigated.
@@ VII-9 Adaptivity
The diverse application scenarios range from transmission to various speech processing tasks and speech LLMs. It is challenging to develop a single speech tokenizer that meets all the requirements of these scenarios. However, adaptivity remains a desirable property, where the tokenizer is expected to adapt to different scenarios given a task prompt or a few data examples. If such tokens exist and the aforementioned challenges are well addressed, then we may be close to achieving a universal speech tokenizer.
@@ VIII Conclusion
Recently, discrete speech tokens have emerged as a rapidly evolving field and a core research direction in the speech LLM era. These tokens encode acoustic or semantic information into a compact discrete representation space, catalyzing the fusion of LLMs and speech processing. Existing discrete speech tokens show rich diversity in model architecture and optimization objectives. In this review, we provide a comprehensive introduction to representative categories of discrete speech tokens, summarizing their motivations and limitations. We conduct a unified analysis of reconstruction and voice conversion across different token types to highlight their unique characteristics. We also review efforts to apply discrete tokens to speech processing tasks, including spoken language understanding, speech generation, and spoken language modeling. Finally, we explore future directions for discrete speech tokenization methods. We hope this review lays a solid foundation for future research in speech technology.